OpenStudy (anonymous):

How do I solve: 6-lx+3l<=-3

OpenStudy (amistre64):

looks like you might wanna subtract 6 from both sides

OpenStudy (amistre64):

then change the signs and flip the inequality for good measure

OpenStudy (anonymous):

Can you draw that for me please? I don't really understand.

OpenStudy (amistre64):

i aint got a good mouse to draw it with; ends up looking like scribble

OpenStudy (anonymous):

Nvm, I got it. The answer is x>=-12 right?

OpenStudy (amistre64):

dunno, lets see: 6-lx+3l<=-3 -6 -6 ------------- -|x+3|<=-9 |x+3| >= 9

OpenStudy (karatechopper):

and then u will have two cases to solve for

OpenStudy (amistre64):

9-3 = 6, and 9+3 = 12 so it looks like x = -12 and 6 for reference points

OpenStudy (karatechopper):

-9>=x+3 OR x+3>=9

OpenStudy (amistre64):

when x=0, 3>= 9 is false so we go out

OpenStudy (karatechopper):

o oops my bad i meant AND

OpenStudy (amistre64):

<--------)--------(---------> -12 6

OpenStudy (amistre64):

x<= -12 union x>=6

OpenStudy (anonymous):

Thank you (:

OpenStudy (amistre64):

hope it helps :)